Working alone, worker A can complete a task in 3 hours; worker B can complete it in 6 hours. How long will it take for both of them, working together, to complete the task?
Answer by eric.schuh(2)

you use this equation 1/a+1/b= 1/t so you will have 1/3+1/6=1/t
3 6 t
you need to find the LCD of 3&6 3*2=6
now it becomes 2t+1t=6 == 3t=6 divide both numbers by 3 3t/3t=t 6/3=2
so t=2 it would take them 2 hours working together